# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land–Surface Predictions

### v1

@@ 4.1 Key Findings and Contributions
As we review the literature, the methods to improve spatial generalization to data‐scarce regions are rare and rarely effective Beery et al. (2018), Gacu et al. (2025). In contrast, StefaLand, combined with lightweight fine‐tuning heads, achieves state-of-the-art or competitive performance across four broad problem classes: streamflow (both CAMELS and global), soil moisture, soil composition while also strengthening the parameterization of differentiable process-based models. Across tasks, the strongest gains come from architectures that fuse StefaLand embeddings with explicit temporal modeling via residual connections, indicating that pretraining on attribute-based spatiotemporal structure yields problem-relevant representations while temporal heads resolve sequence dynamics. These outcomes support the premise that foundation models can democratize prediction quality in data-scarce regions by improving out-of-domain transfer.
The streamflow experiments, spanning basin-scale CAMELS PUB/PUR and global cross-continental runoff, consistently show StefaLand’s improved spatial generalization. The soil moisture results confirm robustness in both random site holdouts and continental transfer (Europe). For soil composition, StefaLand reconciles noisy WoSIS in-situ measurements with improved predictive power over random forests, highlighting its ability to harmonize disparate static datasets.
@@
An additional contribution is computational: our attribute-based approach is markedly more efficient than pixel-wise satellite transformers. The underlying transformer has far fewer parameters (roughly 12 million) and avoids the heavy data management requirements of image-centric pretraining. TerraMind’s larger configuration corresponds to about 7,680 GPU hours Jakubik et al. (2025), Aurora required roughly 14,592 GPU hours Bodnar et al. (2025), and AlphaEarth consumed on the order of 28,672 GPU hours Brown et al. (2025). These figures do not include the petabyte-scale data storage and transfer demands for satellite archives. By comparison, StefaLand’s attribute-based pretraining requires only about 720 GPU hours, making high-quality spatial generalization feasible on budgets accessible to academic groups.
Finally, StefaLand is useful both operationally and as a representation layer. For hydrology, it supports point-scale prediction and regional products, for soil and geomorphic tasks it supplies frozen features that enhance conventional models. This dual role helps bridge a persistent gap between data-rich and data-poor regions by providing transferable geoscience embeddings that downstream models can reliably exploit.
4.2 Limitations
While StefaLand demonstrates strong generalization across diverse tasks, several limitations remain. The pretraining dataset is intentionally narrow and high quality, covering a carefully curated set of forcings and attributes. This focus improves signal extraction but may limit the model’s usefulness for problems outside or only loosely related to those variables, such as deeper subsurface processes like groundwater dynamics. Validation is also biased toward regions with dense observational coverage, such as the United States and Europe; although regional holdout tests suggest good transfer, true performance in data-scarce regions of Africa, Asia, and South America is less certain.
4.3 Future Work
Several directions remain to further develop StefaLand. Expanding the range of targets to include variables such as evapotranspiration, snow water equivalent, and groundwater levels would broaden its applicability, while harmonized global inventories could support more comprehensive evaluation of soil and landslide predictions across differing lithologies and climates. Methodologically, advances such as sequence-aware pretraining, uncertainty-aware prediction heads, and tighter integration with process models offer promising avenues to improve calibration and interpretability while preserving efficiency. Establishing standardized train–test protocols for random and regional holdouts, released with accompanying data splits, will also be important for enabling consistent benchmarking of future geoscience foundation models.
@@ 4.3 Future Work
Overall, StefaLand shows that attribute-centric pretraining combined with lightweight temporal or physics heads can deliver strong spatial generalization across geoscientific tasks while remaining computationally accessible. This points toward a practical path for high-quality predictions in regions where they are most needed but data are most limited.
5 Reproducibility Statement
The pretrained StefaLand model and all code for both pretraining and finetuning is released publicly at [https://anonymous.4open.science/r/StefaLand-9421/]. All datasets used in this work both pretraining dataset and all finetuneing benchmarks are fully public. A complete list of variables used for each task, along with their data sources, is provided in Appendix C. We also report all hyperparameters and model details in the same Appendix.
@@ Appendix A Detailed Model Architecture
This appendix provides the complete mathematical formulation of the StefaLand model architecture.
@@ A.1 Embedding Dynamic and Static Inputs
A.2 Transformer Encoder
@@
A.3 Reconstruction of Original Inputs
@@
Appendix B Physics-Based Differentiable Modeling
To leverage domain knowledge and physical constraints inherent in hydrological systems, we implemented physics-based models that explicitly represent hydrological processes through mathematical formulations. These differentiable versions can be trained end-to-end within neural network frameworks, combining process understanding with machine learning flexibility Shen et al. (2023).
For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model Aghakouchak and Habib (2010), Beck et al. (2020), Bergström (1976; 1992), Seibert and Vis (2012), a relatively simple bucket-type conceptual hydrologic model. HBV has state variables like snow storage, soil water, and subsurface storage, and can simulate flux variables like evapotranspiration (ET), recharge, surface runoff, shallow subsurface flow, and groundwater flow. The parameters of HBV are learned from basin characteristics by a deep learning network.
We used an updated modern version, HBV1.1 Song et al. (2025), which includes modifications such as increased parallel storage components to represent heterogeneity within basins and dynamic parameterization capabilities.
For physics-based configurations, we tested: (1) a baseline LSTM-HBV1.1 configuration as a standard reference, (2) StefaLand HBV1.1 with resConn, which combines the physics-based approach with our residual connection architecture, and (3) StefaLand HBV1.1 without resConn. These physics-based approaches incorporate hydrological process understanding while maintaining the ability to learn from data.
@@ Appendix C Experimental Details
Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP)
Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset
Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3)
Sand, silt, clay fractions Harmonized World Soil Database (HWSD)
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED)
Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers
Carbonate sedimentary rock fraction Global Lithological Map (GLiM)
Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS)
Population density Gridded Population of the World (GPW) v4 dataset
GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015
Forest intact fraction Intact Forest Landscapes Data
@@
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS)
@@
Time Series Forcings Albedo (BSA, WSA) Moderate Resolution Imaging Spectroradiometer (MODIS) MCD43A3 Version 6
LST (Day, Night) MODIS Land Surface Temperature/Emissivity Daily (MYD11A1) Version 6.1
Precipitation Global Precipitation Measurement (GPM), Multi-Source WeightedEnsemble Precipitation (MSWEP)
Forecast albedo, LAI (high/low vegetation), soil temperature (layer 1), surface pressure, solar radiation, 2m temperature, evaporation, precipitation, U/V wind (10m) ERA5-Land Hourly - ECMWF Climate Reanalysis
Static Attributes elevation, slope, aspect, roughness, curvature Global 1,5,10,100-km Topography database
Sand, clay, silt, bulk density Harmonized world soil database (HWSD)
Land cover; urban; open water; snow/ice ESA CCI land cover 2018
NDVI Vegetation Indices Monthly L3 Global 0.05Deg CMG
Outputs Soil moisture ESA CCI SM / in-situ merged
@@
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Radiation, Humidity MSWX and MSWEP
Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset
Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3)
Sand, silt, clay fractions Harmonized World Soil Database (HWSD)
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED)
Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers
Carbonate sedimentary rock fraction Global Lithological Map (GLiM)
Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS)
Population density Gridded Population of the World (GPW) v4 dataset
GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015
Forest intact fraction Intact Forest Landscapes Data
Outputs Streamflow Global gauge networks
@@
Outputs Soil property (clay; sand; silt) The World Soil Information Service (WoSIS)

### v2

@@ 4.1 Key Findings and Contributions
As we review the literature, the methods to improve spatial generalization to data‐scarce regions are rare and rarely effective (Beery et al., 2018, Gacu et al., 2025). In contrast, StefaLand, combined with lightweight fine‐tuning heads, achieves state-of-the-art or competitive performance across four broad problem classes: streamflow (both CAMELS and global), soil moisture, soil composition, and landslide susceptibility, while also strengthening the parameterization of differentiable process-based models. Across tasks, the strongest gains come from architectures that fuse StefaLand embeddings with explicit temporal modeling via residual connections, indicating that pretraining on attribute-based spatiotemporal structure yields problem-relevant representations while temporal heads resolve sequence dynamics. These outcomes support the premise that foundation models can democratize prediction quality in data-scarce regions by improving out-of-domain transfer.
The five dynamic prediction cases, along with benchmark models that reproduce state-of-the-art results in the literature, together paint a clear picture. The pretraining of StefaLand on attributes builds deep landscape representations. These features are highly relevant to hydrologic prediction tasks and can avoid overfitting compared to those built in task-specific supervised learning (LSTM or StefaLand-direct), improving model spatial generalization. They are also decidedly more relevant to such tasks than those obtained from existing Earth Foundation models expensively trained on massive amounts of satellite images. It shows that the importance of larger pretraining data may not necessarily exceed that of problem relevance, and image-like data may not be the optimal data representation for such tasks. We stress that our approach is complementary to satellite-based foundation models: where they exploit large-scale visual patterns, StefaLand focuses on problem-relevant attributes, offering an efficient and domain-specific alternative. We have tested a large number of adapter formulations. While we can certainly test more finetuning options, especially internal foundation model layers, the computational and data-storage demands already start to be limiting.
@@
Our attribute-based approach is at least an order of magnitude more efficient than pixel-wise satellite transformers. Our underlying transformer has far fewer parameters (roughly 12 million) and avoids the heavy data management requirements of image-centric pretraining. TerraMind’s larger configuration corresponds to about 7,680 GPU hours (Jakubik et al., 2025), PrithviWxC require roughly 23,040 GPU hours (Schmude et al., 2024) and Aurora required roughly 14,592 GPU hours (Bodnar et al., 2025). In contrast, StefaLand’s attribute-based pretraining requires only about 720 GPU hours, making high-quality spatial generalization feasible on limited budgets. These figures do not include the large gaps in data-storage and transfer demands during pretraining (StefaLand, Terramind and PrithviWxC used $\approx 2$, $\approx 11$ and $\approx 27$ data as mentioned earlier). In fact, it is costly and resource-straining to run the satellite-imagery-focused foundation models, which impedes us from comparing with them in every case. Finally, to the best of our effort, we could not identify other earth foundation models that are designed for land-surface predictions.
4.2 Limitations and Future Work
@@
Several limitations remain. The selection of geological- or ecologically-focused attributes is limited and more can be added to further characterize the subsurface. Two-dimensional (or image-like) data like elevation map can be selectively incorporated using vision transformer heads in the future. Expanding the range of targets to include variables such as evapotranspiration, snow water equivalent, and groundwater levels would broaden its applicability. Methodologically, advances such as uncertainty-aware prediction heads, and tighter integration with additional process models offer promising avenues to improve calibration and interpretability while preserving efficiency. Overall, StefaLand shows that attribute-centric pretraining combined with lightweight temporal or physics heads can deliver strong spatial generalization across geoscientific tasks while remaining computationally accessible. This points toward a practical path for high-quality predictions in regions where they are most needed but data are most limited.
@@ Appendix A Detailed Model Architecture
This appendix provides the complete mathematical formulation of the StefaLand model architecture more detailed then 2.
@@ A.1 Embedding Dynamic and Static Inputs
A.2 Location-aware Cross-Variable Group Masking
StefaLand introduces Cross-Variable Group Masking (CVGM), a masking strategy that forces the model to capture interactions among correlated hydrologic variables rather than treating them independently. Given a predefinition of hydrological variables into groups $G=\{g_{1},g_{2},\ldots,g_{k}\}$, masking occurs as follows:
  1. 1.
    A temporal masking window $\tau,\tau+\ell)$ is randomly sampled with length $\ell\sim\mathcal{U}(L_{\text{min}},L_{\text{max}})$.
  2. 2.
    For each variable group $g_{k}$, a Bernoulli mask indicator $m_{k}\sim\text{Bernoulli}(p_{\text{mask}})$ determines whether the group is masked.
  3. 3.
    For each time step $t$ within the masked temporal window and each variable $c$ belonging to masked groups, the embedded feature vector is replaced by a learned mask vector $\mathbf{m}_{c}$.
Each hydrologic variable $c$ has its own trainable mask embedding vector $\mathbf{m}_{c}\in\mathbb{R}^{256}$. This CVGM procedure creates reconstruction targets that require modeling cross-variable dependencies and physical interactions.
A.3 Learnable Positional Encoding
To provide positional information, StefaLand employs learnable positional encoding. Each position $i$, corresponding to each time step and the appended static embedding, is assigned a trainable embedding vector $\mathbf{p}_{i}$. The encoded embedding becomes:
$\tilde{Z}=Z+P$ (20)
where $P=[\mathbf{p}_{1};\ldots;\mathbf{p}_{T+1}]$.
A.4 Transformer Encoder
@@
A.5 Reconstruction of Original Inputs
@@
Appendix B Additional Experiments
B.1 Physics-Based Differentiable Modeling
To leverage domain knowledge and physical constraints inherent in hydrological systems, we implemented physics-based models that explicitly represent hydrological processes through mathematical formulations. These differentiable versions can be trained end-to-end within neural network frameworks, combining process understanding with machine learning flexibility (Shen et al., 2023).
For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model (Aghakouchak and Habib, 2010, Beck et al., 2020, Bergström, 1976; 1992, Seibert and Vis, 2012), a relatively simple bucket-type conceptual hydrologic model. HBV has state variables like snow storage, soil water, and subsurface storage, and can simulate flux variables such as evapotranspiration (ET), recharge, surface runoff, shallow subsurface flow, and groundwater flow. We used an updated modern version, HBV1.1 (Song et al., 2025), which includes modifications such as increased parallel storage components to represent heterogeneity within basins and dynamic parameterization capabilities.
The hybrid model employs a differentiable parameter learning (dPL) framework where neural networks generate parameters for HBV1.1, and errors are backpropagated through the entire system. A machine learning network takes basin attributes and meteorological forcings as inputs and outputs HBV parameters—both static (e.g., recession coefficients) and dynamic parameters that vary daily. Because HBV1.1 supports automatic differentiation, it serves as the physical backbone: during training, loss is calculated between simulated and observed streamflow, gradients are backpropagated through HBV equations, and neural network weights are updated. This differs from traditional calibration because parameters are learned regionally across all basins simultaneously rather than individually, allowing the network to capture generalizable relationships between basin characteristics and optimal parameters while maintaining mass balance constraints. The system uses 16 parallel response units for spatial heterogeneity and outputs diagnostic variables (e.g., evapotranspiration, soil moisture, baseflow) not directly trained on, providing interpretability with competitive performance.
For physics-based configurations, we tested: (1) a baseline LSTM–HBV1.1 configuration as a standard reference, (2) StefaLand HBV1.1 with resConn, which combines the physics-based approach with our residual connection architecture, and (3) StefaLand HBV1.1 without resConn. These physics-based approaches incorporate hydrological process understanding while maintaining the ability to learn from data..
Table B1: CAMELS Streamflow PUB and PUR Results (Physics-Based Models)
RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
LSTM - HBV1.1 1.325 1.298 0.857 0.672 1.561 1.521 0.746 0.578
B.2 Linear Regression baselines
To justify the use of complex neural networks over traditional methods, we have conducted baseline comparisons using linear regression models. As shown in the table below, linear regression performs poorly across all tasks by a fair margin when compared to our neural network approaches.
Table B2: Additional experiments with linear regression baselines.
Experiment Random holdout Regional holdout
RMSE ↓ µbRMSE ↓ Corr ↑ RMSE ↓ µbRMSE ↓ Corr ↑
Soil Moisture Linear Regression 0.120 0.101 0.188 0.121 0.103 0.187
@@ Appendix C Experimental Details
Table C1: StefaLand Pretraining Configuration
Parameter Value
General Settings
Task pretrain
Model MFFormer_dec_LSTM
Random seed 111
Time Period 1980/1/1–2018/12/31
Sequence Configuration
Sequence length 365
Label length 365
Prediction length 365
Sampling stride 1
Minimum window size 30
Maximum window size 90
Model Architecture
Input dimension (enc_in) 32
Decoder input (dec_in) 6
Output dimension (c_out) 6
Model dimension 256
Number of heads 4
Encoder layers 4
Decoder layers 2
Feed-forward dimension 512
Dropout 0.1
Activation gelu
Training Configuration
Optimizer AdamW
Loss criterion MaskedNSE
Epochs 25
Batch size 256
Learning rate 0.0001
Weight decay 0.0
Patience 30
Gradient clipping 5.0
Number of workers 10
Loss Weights
Time series loss ratio 1.0
Static loss ratio 0.5
@@
Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP) (Beck et al., 2022; 2019)
Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset (ESA, 2017)
Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3) (Didan, 2015a)
Sand, silt, clay fractions Harmonized World Soil Database (HWSD) (FAO et al., 2012)
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011, Ramcharan et al., 2018)
Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers (Pelletier et al., 2016)
Carbonate sedimentary rock fraction Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012)
Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS) (Gleeson et al., 2014)
Population density Gridded Population of the World (GPW) v4 dataset (CIESIN, 2016)
GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015 (Kummu et al., 2018)
Forest intact fraction Intact Forest Landscapes Data (Potapov et al., 2017)
@@
Table C3: Attribute Groups Used in Group Masking Pretraining
Group Variables
Topography meanelevation, meanslope
Soil HWSD_clay, HWSD_sand, HWSD_silt, HWSD_gravel, SoilGrids1km_sand, SoilGrids1km_clay, SoilGrids1km_silt
Geology permeability, Porosity, glaciers, permafrost
Vegetation NDVI, FW
Climate aridity, meanP, ETPOT_Hargr, meanTa, seasonality_P, seasonality_PET, snow_fraction, snowfall_fraction
@@
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS) (Addor et al., 2017, Newman et al., 2014)
@@
Time Series Forcings Albedo (BSA, WSA) Moderate Resolution Imaging Spectroradiometer (MODIS) MCD43A3 version 6 (Schaaf, Crystal and Wang, Zhuosen, 2021)
LST (Day, Night) MODIS Land Surface Temperature/Emissivity Daily (MYD11A1) Version 6.1 (Wan et al., 2021)
Precipitation Global Precipitation Measurement (GPM), & MSWEP &ERA5 precipitation (Huffman et al., 2019, Beck et al., 2019, Muñoz Sabater, 2019)
Forecast albedo, LAI (high/low vegetation), soil temperature (layer 1), surface pressure, solar radiation, 2 m temperature, evaporation, precipitation, U/V wind (10 m) ECMWF Reanalysis v5 (ERA5) (Muñoz Sabater, 2019)
Static Attributes elevation, slope, aspect, roughness, curvature Global 1/5/10/100-km topography derivatives (Amatulli et al., 2018)
Sand, clay, silt, bulk density HWSD v1.2 (FAO et al., 2012)
Land cover; urban; open water; snow/ice ESA CCI Land Cover (ESA, 2017)
NDVI Vegetation Indices Monthly L3 Global 0.05Deg CMG (Didan et al., 2015)
Outputs Soil moisture International Soil Moisture Network (ISMN) (Dorigo et al., 2013b; 2011)
@@
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Radiation, Humidity MSWX and MSWEP (Beck et al., 2022; 2019)
Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset (ESA, 2017)
Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3) (Didan, 2015a)
Sand, silt, clay fractions Harmonized World Soil Database (HWSD) (FAO et al., 2012)
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011, Ramcharan et al., 2018)
Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers (Pelletier et al., 2016)
Carbonate sedimentary rock fraction Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012)
Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS) (Gleeson et al., 2014)
Population density Gridded Population of the World (GPW) v4 dataset (CIESIN, 2016)
GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015 (Kummu et al., 2018)
Forest intact fraction Intact Forest Landscapes Data (Potapov et al., 2017)
Outputs Streamflow Global runoff data (GRDC, 2024)
Table C9: Landslide (SLIDO, Oregon) Variables and Sources
Variable Type Variable Name Source
Input data Elevation National Elevation Dataset (NED) (Gesch et al., 2018)
Soil sand, silt, clay, bulk density, saturated hydraulic conductivity Probabilistic Remapping of SSURGO (POLARIS) (Chaney et al., 2019)
Lithology Global Lithological Map (GLiM) (har)
Rainfall PRISM (PRISM Climate Group, 2014)
NDVI Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices Monthly L3 (Didan, 2015b)
Landcover National Land Cover Database (NLCD) 2016 (Dewitz, 2019)
Soil moisture SMAP-HydroBlocks (SMAP-HB) (Vergopolan et al., 2021)
slope, aspect, curvature, TWI, SPI DEM-derived
Outputs Landslide occurrence (binary) Statewide Landslide Information Database for Oregon (SLIDO) (Franczyk, J. J et al., 2020)
@@
Outputs Soil property (clay; sand; silt) World Soil Information Service (WoSIS) (Batjes et al., 2020)
Table C11: PrithviWxC Surface Variables and Sources (MERRA-2)
Variable Type Variable Name Source
Time Series Forcings Precipitation, fluxes, winds, soil moisture/temperature, LAI, runoff, etc. NASA MERRA-2 reanalysis (Gelaro et al., 2017)
Static Attributes Land/ocean/ice fractions, surface geopotential, subgrid orography NASA MERRA-2 constants (Gelaro et al., 2017)
@@
C.1 Pretraining Data Handling
To handle large volumes efficiently, data were stored in shards rather than fully loaded into memory. At the start of each epoch, shards were randomly reindexed, providing diverse samples while avoiding costly materialization of the full dataset. This strategy ensures both reproducibility and wide sample coverage. The accompanying code release in the reproducibility statement enables reconstruction of this dataset and reproduces the full pretraining pipeline.
C.2 Dataset Splitting
For the WoSIS soil dataset, we collected soil property data from 106,503 locations. After removing low-quality records (e.g., sand values greater than 1 or negative values), we randomly sampled 5,000 soil points to reduce computational cost. We then applied 5-fold cross-validation (k=5) on this subset.
For the landslide dataset at 30 m resolution, we used 14,604 historical landslide points. We split the dataset into 70% for training, 20% for validation, and 10% for testing.